OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = -10 3x - 3y + 6z = -21 x + 2y + 2z = -10 Can somebody do this out and show me how it's done? Thanks :]

OpenStudy (anonymous):

sure...

OpenStudy (anonymous):

I think that if you combine the equations in such a way as to eliminate variables, u can then solve for one of them. Then you can plug that number back in and solve for the rest!

OpenStudy (anonymous):

u may apply elimination method

OpenStudy (anonymous):

see, u hv three variables -- x, y and z and so three equations x + 4y + z = -10 ----------------(1) 3x - 3y + 6z = -21 ----------------(2) x + 2y + 2z = -10 ----------------(3) We have to use these equations in pairs to eliminate two of the variables so that only one is left. This we do as follows: First we take (1) and (2) x + 4y + z = -10 ----------------(1) 3x - 3y + 6z = -21 ----------------(2) Now multiply (1) by 3 to get equation (4) and subtract (2) from it 3x + 12y + 3z = -30 ----------------(4) 3x - 3y + 6z = -21 ----------------(2) we get 15y - 3z = -9 ----------------(A) Then we take (2) and (3) 3x - 3y + 6z = -21 ----------------(2) x + 2y + 2z = -10 ----------------(3) Now we multiply (3) by 3 to get eqn (5) and subtract it from (2) 3x - 3y + 6z = -21 ----------------(2) 3x + 6y + 6z = -30 ----------------(5) we get -9y = 9 y = 9/-9 y = -1

OpenStudy (anonymous):

understood till this point.....????

OpenStudy (anonymous):

Yep, I understand :]

OpenStudy (anonymous):

Now substitute the y = -1 in equation (A) 15(-1) - 3z = -9 -15 -3z = -9 -3z = -9 + 15 z = 6/-3 =-2 z = -2 so till now we have y = -1 and z = -2 got it...????

OpenStudy (anonymous):

Now substitute these values of y and z in (1) x + 4y + z = -10 x + 4(-1) + (-2) = -10 x - 4 - 2 = -10 x = -10 + 6 x = -4 so solutions set (x,y,z) = (-4, -1, -2)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

u r welcome.... ☺